Find an equation of the line described. Write the equation in the slope-intercept formif possible.
Through the origin and (6, 3/2)

Answer by richwmiller(17219)

so you have to points (0,0) and (6,3/2)
y-y=m(x-x)
where m is the slope
m=3/2-0/6-0=3/2/6
multiply both num and den by 2
m=3/12
y=mx+b
0=0+b
b=0
y=3/12*x
y=3x/12